# Real-time operator evolution in two and three dimensions via sparse Pauli dynamics

### v1

@@ II Real-time sparse Pauli dynamics
III Results
III.1 Spin and energy diffusion constants in 1D chains
@@ III.1.1 Models
III.1.2 Numerical results
Figure 1 shows the mean-square displacements for the tilted-field Ising (panel A) and XX-ladder models (panel C), simulated using SPD with different values of the threshold $\delta$ ranging from $2^{-18}$ to $2^{-13}$. We can observe how $d^{2}(t)$ converges onto a straight line at large $t$ as we reduce the threshold and make the simulations more accurate. Figures 1B, D plot the corresponding diffusion constant as a function $\delta/\Delta t$ for two sets of data, one using $\Delta t=0.01$ and the other using $\Delta t=0.02$. For this property, the results are amenable to an extrapolation in $\delta\to 0$, and we find $D\approx 1.4$ for the tilted-field Ising chain (same value was reported in Ref. [4]), while $D\approx 0.94$ for the XX-ladder (compared to $D\approx 0.95$ of Ref. [30] and $D\approx 0.96-0.98$ reported in Ref. [4]). The plots of the diffusion constant also reveal the scaling relationship discussed in Sec. II. Namely, as the threshold is reduced, the simulations using different $\delta$ and $\Delta t$ but the same $\delta/\Delta t$ become closer to each other. In addition, we can numerically verify that using a larger time step reduces the computational cost. For example, the two points in Fig. 1D at fixed $\delta/\Delta t=2^{-18}/0.02\approx 0.00019$ take around 848484\,84h ($\Delta t=0.01$, blue circle) and 434343\,43h ($\Delta t=0.02$, orange square) to simulate on 6 cores.
Figure 1: Time-dependent mean-square displacements (A, C) and diffusion constants (B, D) for the tilted-field Ising (A, B) and XX-ladder (C, D) models. Extrapolations correspond to linear fits to the leftmost three points.
@@ III.1.2 Numerical results
Figure 2: Comparison of SPD and MPO on the example of short-chain ($L=9-21$) tilted-field Ising model. Exact benchmarks were adapted from Ref. [4].
@@
Figure 3: A: Contributions of Paulis of different weights $m$ to the total Frobenius norm squared $F$ of the operator $q_{\frac{L+1}{2}}(t)$ [Eq. (12)] evolved under the tilted field Ising Hamiltonian [Eq. (11)]. B: Individual Pauli weight contributions $F_{m}$ exhibit expected asymptotic behavior of $t^{-1/2}$ for conserved properties (local energy that consists of $m=1$ and $m=2$ Paulis) and $t^{-3/2}$ for other Pauli operators ($m>2$). In these examples, the simulations were performed up to $t=40$ with a threshold of $2^{-17}$. All data were rescaled by $F(0)=\text{Tr}[q_{\frac{L+1}{2}}(0)^{2}]/2^{n}=3.27812025$.
@@
III.2 Time-dependent expectation values
III.2.1 X-truncated sparse Pauli dynamics
@@
III.2.2 2D transverse-field Ising model
@@
Figure 4: Convergence of xSPD ($M=5$) with respect to threshold and comparison with the iPEPS simulations based on full update (FU) [38], neighborhood tensor update (NTU) [39], and gradient tensor update (GTU) [40] schemes. The plot in the right panel zooms in on the last part of dynamics ($t>0.7$). All data are plotted at intervals of $\Delta t=0.04$.
@@
Figure 5: A: We estimate the time step error by comparing different time steps at a fixed $\delta/\Delta t$ ratio. We find that the Trotter error for $\Delta t=0.04$ is less than 0.005. B: The error of using xSPD, for truncation with $M=3,5,7$. We find that $M=5$ is optimal choice for the total time and accuracy we target in our simulation, with an error below 0.005.
@@
III.2.3 3D transverse-field Ising model
@@
The system with $h=5.158136$ (Fig. 6B) was propagated up to $t=0.6$ using $\Delta t=0.02$ and $M=5$. For this choice of parameters, the Trotter and X-truncation errors are estimated to be below 0.005 (see Fig. 8B, D in Appendix C). Because of the shorter dynamics we could use a smaller value of the X-truncation parameter $M$ compared to the weak-field case. The results, using thresholds as low as $\delta=2^{-20}$, are converged to below 0.01 for times $t<0.5$, after which our most accurate simulations begin to deviate from each other.
@@
Figure 6: Convergence of xSPD with respect to threshold for the 3D system at $h=1$ (panel A, using $\Delta t=0.04$ and $M=7$) and $h=5.158136$ (panel B, using $\Delta t=0.02$ and $M=5$). Inset in panel A shows error with respect to the most accurate available simulation ($\delta=2^{-19}$).
III.2.4 Computational savings from X-truncation
@@
Figure 7: The effect of X-truncation on the number of Pauli operators. A, B: Number of all (black) and Z-type (red) Pauli operators generated with SPD (solid) or xSPD (dashed) for the 2D square lattice (A, threshold $\delta=2^{-20}$) and 3D cubic lattice (B, $h=1$, threshold $\delta=2^{-16}$). C: Ratio between the number of Pauli operators in the observable of SPD and xSPD as a function of time.
IV Conclusion
@@ Appendix A Additional details of SPD implementation and working equations
In Sec. III.1, we also introduced an overlap (inner product) between sums of Pauli operators $\text{Tr}[O_{1}^{\dagger}O_{2}]/2^{n}$ represented in the sparse Pauli format described above. Let us assume that $N_{1}<N_{2}$, i.e., that $O_{1}$ has fewer Pauli operators than $O_{2}$. Then, we can search for all Pauli operators of $O_{1}$ in $O_{2}$ in $\mathcal{O}(N_{1}\log N_{2})$ time and the overlap is
@@
where the sum runs over Pauli operators in $O_{1}$ that were found in $O_{2}$ and $k[j]$ is the index of the $j$-th found Pauli in $O_{2}$. The expectation over the all-zero state, as needed in Sec. III.2, can be computed as

### v2

@@ II Real-time sparse Pauli dynamics
Finally, we introduce an upper bound on the number of Pauli operators ($N$) for a given threshold $\delta$. Let us assume that the scaled Frobenius norm squared of the observable operator $O_{0}$ before propagation is
$\frac{1}{2^{n}}\text{Tr}(O_{0}^{\dagger}O_{0})=c_{0}.$ (8)
Then, during propagation, $O_{t}=\sum_{P}a_{P}P$, and this norm is preserved under unitary dynamics but is otherwise smaller than $c_{0}$ if we truncate the observable based on sparse Pauli dynamics. Therefore, we have
$\frac{1}{2^{n}}\text{Tr}(O_{t}^{\dagger}O_{t})=\sum_{P}|a_{P}|^{2}\leq c_{0}.$ (9)
Because of the truncation based on the threshold $\delta$, we know that
$\sum_{P}|a_{P}|^{2}\geq\sum_{P}\delta^{2}=N\delta^{2},$ (10)
which leads to
$N\leq\frac{c_{0}}{\delta^{2}}.$ (11)
III Related work and theoretical guarantees
Sparse Pauli dynamics was developed in the context of quantum circuit simulation and originates from the idea that the sparse Pauli representation can be efficiently updated under Clifford operations [23, 17, 18]. In turn, for a fixed threshold, it typically exhibits an exponential scaling of the number of Pauli operators with respect to the number of non-Clifford gates. One way to analyze methods based on the sparse Pauli representation is by considering magic monotones, such as the recently developed operator stabilizer entropy [28]. This magic monotone, which has been shown to satisfy the Lieb-Robinson bound [29], is closely related to the number of Pauli operators one must store to retain a faithful representation of the observable operator.
One limit in which sparse Pauli dynamics is guaranteed to perform well is when the dynamics is constrained to a subset of Pauli operators that is much smaller than the operator Hilbert space (e.g., of polynomial size in $n$). This happens for certain observable evolutions in integrable systems, such as in the non-interacting transverse-field Ising chain. In Fig. 1, we show an example of such dynamics where the number of Pauli operators is limited. For the studied example, the Pauli operator $Z$ maps to a two-index Majorana operator and the Hamiltonian maps to a free fermion model, implying that the number of Paulis in the sparse Pauli representation saturates at $\binom{2n}{2}$ [30, 31]. Finite-threshold sparse Pauli dynamics still provides a significant computational advantage over exact simulation due to the fact that many Paulis do not contribute to the observable at early times (compare the $\delta=0$ and $\delta\neq 0$ results in Fig. 1). The above mentioned operator stabilizer entropy has also been derived for the dual-unitary XXZ model, a type of local interacting integrable circuit, where it was shown to saturate to a constant at long times [28], thus providing another analytical example where the sparse Pauli representation can be shown to be asymptotically efficient.
Figure 1: An example of observable dynamics in an integrable system for which the observable is limited to a polynomial number of Pauli operators in system size $n$. We compute $\langle Z_{0}\rangle$, where the zero index indicates the central site of a length $n=101$ chain, time evolved with the transverse-field Ising model $H=\sum_{i}X_{i}X_{i+1}+Z_{i}$, starting from an all-zero state $|0^{\otimes n}\rangle$. The dynamics is simulated for $t=40$ with $\Delta t=0.02$ and different values of the threshold $\delta$. Solid lines correspond to the expectation values (left axis), while the dashed lines correspond to the number of Pauli operators at each time step (right axis). The horizontal gray line is the maximum number of Pauli operators $\binom{2n}{2}=20301$ that can be generated in this case.
Finally, Pauli-based techniques have been of central interest for analyzing the classical simulability of noisy quantum circuits, including the computation of probability amplitudes and expectation values of Pauli operators. Specifically, it has been shown that they enable simulation of noisy quantum circuits with a cost that is exponential only in the inverse of the noise rate but polynomial in the size of the system and depth of the circuit [21, 16, 20, 22]. More recently, these results have been generalized to noiseless, locally scrambling circuits [32]. Sparse Pauli dynamics profits directly from these theoretical results even though the theoretical analysis is based primarily on weight-based Pauli truncation, as there is a strong correlation between high weight Paulis and small coefficients.
However, unlike the above cases that come with some theoretical guarantees, general quantum dynamics comes with no guarantees and is expected to require computational resources that scale exponentially in time. This is shared by all classical numerical methods for simulating quantum dynamics. Therefore, in this setting we can analyze the merits of individual methods only empirically. In remainder of this paper we therefore focus on an empirical analysis of the numerical results we obtain for real-time quantum dynamics.
IV Results
IV.1 Spin and energy diffusion constants in 1D chains
@@ IV.1.1 Models
IV.1.2 Numerical results
Figure 2 shows the mean-square displacements for the tilted-field Ising (panel A) and XX-ladder models (panel C), simulated using SPD with different values of the threshold $\delta$ ranging from $2^{-18}$ to $2^{-13}$. We can observe how $d^{2}(t)$ converges onto a straight line at large $t$ as we reduce the threshold and make the simulations more accurate. Figures 2B, D plot the corresponding diffusion constant as a function $\delta/\Delta t$ for two sets of data, one using $\Delta t=0.01$ and the other using $\Delta t=0.02$. For this property, the results are amenable to an extrapolation in $\delta\to 0$, and we find $D\approx 1.4$ for the tilted-field Ising chain (same value was reported in Ref. [4], whereas somewhat larger values of $D\approx 1.44$ [38] and $D\approx 1.55$ [39] were found in more recent works), while $D\approx 0.94$ for the XX-ladder (compared to $D\approx 0.95$ of Refs. [35, 37, 40] and $D\approx 0.96-0.98$ reported in Ref. [4]). The plots of the diffusion constant also reveal the scaling relationship discussed in Sec. II. Namely, as the threshold is reduced, the simulations using different $\delta$ and $\Delta t$ but the same $\delta/\Delta t$ become closer to each other. In addition, we can numerically verify that using a larger time step reduces the computational cost. For example, the two points in Fig. 2D at fixed $\delta/\Delta t=2^{-18}/0.02\approx 0.00019$ take around 848484\,84h ($\Delta t=0.01$, blue circle) and 434343\,43h ($\Delta t=0.02$, orange square) to simulate on 6 cores.
Figure 2: Time-dependent mean-square displacements (A, C) and diffusion constants (B, D) for the tilted-field Ising (A, B) and XX-ladder (C, D) models described in Sec. IV.1.1. Extrapolations correspond to linear fits to the leftmost three points.
@@ IV.1.2 Numerical results
Figure 3: Comparison of SPD and MPO on the example of short-chain ($L=9-21$) tilted-field Ising model of Sec. IV.1.1. Exact benchmarks were adapted from Ref. [4]. Typical runtimes of MPO and SPD simulations are given in Table 1, Appendix B.
@@
Figure 4: A: Contributions of Paulis of different weights $m$ to the total Frobenius norm squared $F$ of the operator $q_{\frac{L+1}{2}}(t)$ [Eq. (16)] evolved under the tilted field Ising Hamiltonian [Eq. (15)] (see Sec. IV.1.1). B: Individual Pauli weight contributions $F_{m}$ exhibit expected asymptotic behavior of $t^{-1/2}$ for conserved properties (local energy that consists of $m=1$ and $m=2$ Paulis) and $t^{-3/2}$ for other Pauli operators ($m>2$). In these examples, the simulations were performed up to $t=40$ with a threshold of $2^{-17}$. All data were rescaled by $F(0)=\text{Tr}[q_{\frac{L+1}{2}}(0)^{2}]/2^{n}=3.27812025$.
@@
IV.2 Time-dependent expectation values
IV.2.1 X-truncated sparse Pauli dynamics
@@
IV.2.2 2D transverse-field Ising model
@@
Figure 5: Convergence of xSPD ($M=5$) with respect to threshold and comparison with the iPEPS simulations based on full update (FU) [46], neighborhood tensor update (NTU) [47], and gradient tensor update (GTU) [48] schemes for the $11\times 1$ 11 2D square-lattice transverse-field Ising model of Eq. (18) at critical point $h=h_{c}=3.044382$. The plot in the right panel zooms in on the last part of dynamics ($t>0.7$). All data are plotted at intervals of $\Delta t=0.04$.
@@
Figure 6: Time step and X-truncation errors for the $11\times 1$ 11 2D square lattice model at critical point $h=h_{c}=3.044382$. A: We estimate the time step error by comparing different time steps at a fixed $\delta/\Delta t$ ratio. We find that the Trotter error for $\Delta t=0.04$ is less than 0.005. B: The error of using xSPD, for truncation with $M=3,5,7$. We find that $M=5$ is optimal choice for the total time and accuracy we target in our simulation, with an error below 0.005.
@@
IV.2.3 3D transverse-field Ising model
@@
The system with $h=5.158136$ (Fig. 7B) was propagated up to $t=0.6$ using $\Delta t=0.02$ and $M=5$. For this choice of parameters, the Trotter and X-truncation errors are estimated to be below 0.005 (see Fig. 9B, D in Appendix C). Because of the shorter dynamics we could use a smaller value of the X-truncation parameter $M$ compared to the weak-field case. The results, using thresholds as low as $\delta=2^{-20}$, are converged to below 0.01 for times $t<0.5$, after which our most accurate simulations begin to deviate from each other. This example also illustrates one of the limitations of SPD when it comes to converging results with respect to threshold. Namely, the expectation values are not monotonically converging to the exact result, implying that it is difficult to perform zero-threshold extrapolation, at least far from convergence. Nonetheless, as we approach converged results, such as at smaller thresholds in the 2D square lattice example (Fig. 5), we begin to observe systematic improvement as we reduce the threshold.
@@
Figure 7: Convergence of xSPD with respect to threshold for the $11\times 11\times 1$ 11 3D tranverse-field Ising model at $h=1$ (panel A, using $\Delta t=0.04$ and $M=7$) and $h=5.158136$ (panel B, using $\Delta t=0.02$ and $M=5$). Inset in panel A shows error with respect to the most accurate available simulation ($\delta=2^{-19}$).
IV.2.4 Computational savings from X-truncation
@@
Figure 8: The effect of X-truncation on the number of Pauli operators. A, B: Number of all (black) and Z-type (red) Pauli operators generated with SPD (solid) or xSPD (dashed) for the 2D square lattice (A, $h=3.044382$, threshold $\delta=2^{-20}$) and 3D cubic lattice (B, $h=1$, threshold $\delta=2^{-16}$). C: Ratio between the number of Pauli operators in the observable of SPD and xSPD as a function of time.
V Conclusion
@@ Acknowledgements.
Data availability
The data that support the findings of this article are openly available in a Zenodo repository [51]. The code used to produce the results is available at https://github.com/tbegusic/spd.
@@ Appendix A Additional details of SPD implementation and working equations
In Sec. IV.1, we also introduced an overlap (inner product) between sums of Pauli operators $\text{Tr}[O_{1}^{\dagger}O_{2}]/2^{n}$ represented in the sparse Pauli format described above. Let us assume that $N_{1}<N_{2}$, i.e., that $O_{1}$ has fewer Pauli operators than $O_{2}$. Then, we can search for all Pauli operators of $O_{1}$ in $O_{2}$ in $\mathcal{O}(N_{1}\log N_{2})$ time and the overlap is
@@
where the sum runs over Pauli operators in $O_{1}$ that were found in $O_{2}$ and $k[j]$ is the index of the $j$-th found Pauli in $O_{2}$. The expectation over the all-zero state, as needed in Sec. IV.2, can be computed as
@@ Appendix B Details of MPO simulations
Table 1: Runtimes of MPO and SPD simulations for the 1D tilted-field Ising model dynamics shown in Fig. 3, $L=21$. Calculations were performed on 6 cores of an Intel Xeon Platinum 8352Y processor (2.2 GHz).
MPO SPD
$\chi$ Runtime (minutes) $\delta$ Runtime (minutes)
$2^{6}$ 14 $2^{-13}$ 5
$2^{7}$ 44 $2^{-14}$ 21
$2^{8}$ 256 $2^{-15}$ 89
$2^{9}$ 1901 $2^{-16}$ 453